# Hidden Bose-Einstein Singularities of Correlated Electron Systems: III. Thermodynamic Signals

### v1

@@ I Introduction
Quantum field theory has proved extremely useful for describing properties of many-particle systems, as can be seen by looking into clalssic textbooks.KB62 ; AGD63 ; FW71 ; NO88 ; Mahan00 However, the main focus there is on Green’s functions that are relevant to one- and two-particle properties rather than thermodynamic potentials. Moreover, ordered phases such as superconductivity are treated almost exclusively within the mean-field framework. Indeed, there seems to have been few detailed studies on how to calculate thermodynamic potentials of normal and ordered phases with correlations microscopically and continuously through phase transitions. One of the exceptional ones was given by Rainer and SereneRS76 ; SR83 on thermodynamics of superfluid 3He, where correlation effects are treated perturbatively as strong-coupling effects on the mean-field results by using the LW functional. Another study to be mentioned was given by Haussmann et alHRCZ07 on thermodynamics of the BCS-BEC crossover of atomic gases based on the attractive Hubbard model, who adopted the De Dominicis-Martin (DDM) formalismDDM64 of expressing $\Omega$ as a functional of Green’s functions and two-particle vertices in addition to the LW framework. However, (i) their mixed use of the DDM and LW formalisms and (ii) the renormalization of the bare interaction vertex obscure the exact procedure to obtain the thermodynamic potentials.
With these backgrounds, the aims of the present paper are twofold. First, we establish how to calculate thermodynamic potentials with correlations numerically with high precision based on the LW functional alone. Using them, we perform a numerical study on the Helmholtz free energy ${\cal F}$, internal energy ${\cal E}$, entropy ${\cal S}$, and heat capacity ${\cal C}_{V}$ of the weakly attractive Hubbard model investigated previouslyKita24 ; Kita25 to clarify how its hidden BE singularity is detectable thermodynamically. It will be shown that the entry into the pseudogap phase with the 1PR structure in the self-energy is distinguishable by cusps in ${\cal E}$ and ${\cal S}$ and discontinuity in ${\cal C}_{V}$, similarly as those of the mean-field second-order transitions with broken symmetries.
@@ II.1.1 Grand thermodynamic potential
×$\,\times\frac{\delta\Sigma(\vec{k}^{\prime})}{\delta G(\vec{k})}e^{i\varepsilon_{n^{\prime}}0_{+}}-2T\Sigma(\vec{k})+\frac{\delta\Phi}{\delta G(\vec{k})}.$ (11)
@@ II.1.2 $\Phi$ functional
which holds for the second-order contribution to $\Phi$ in the normal state. However, the former expression with $\chi_{\rm GG}(\vec{q})$ is far more suitable numerically because of the inequality $|\chi_{GG}(\vec{q})|\ll|\chi_{G\bar{G}}(\vec{q})|$ that holds for the attractive Hubbard model at low temperatures. Indeed, Eq. (17) indicates that the tiny second-order contribution given in terms of the particle-hole bubble $\chi_{GG}(\vec{q})$ results by the expression in terms of the particle-particle bubble $\chi_{G\bar{G}}(\vec{q})$ through the cancelation of its huge real and imaginary numbers. Hence, we have calculated the normal-state $\Phi$ by Eq. (14) above.
@@ II.1.3 Other thermodynamic potentials
The Helmholtz free energy ${\cal F}$ is obtained from Eq. (10) by the thermodynamic relation
@@
where ${\cal N}_{\rm e}$ is the electron number defined by Eq. (8). Second, the internal energy ${\cal E}$ can be calculated by the expression
@@
to obtain ${\cal E}$ from Eq. (18) by numerical differentiation. Third, entropy ${\cal S}$ can be calculated by either of the thermodynamic relations
@@ II.2.2 $\Phi$ functional
$\underline{\chi}^{({\rm c})}(\vec{q}\,)\equiv\begin{bmatrix}\vspace{1mm}\chi_{-}(\vec{q}\,)&\sqrt{2}\chi_{GF}(\vec{q}\,)&-\sqrt{2}\chi_{\bar{G}F}(\vec{q}\,)\\ \vspace{1mm}\sqrt{2}\chi_{GF}(\vec{q}\,)&-\chi_{G\bar{G}}(\vec{q}\,)&-\chi_{FF}(\vec{q}\,)\\ -\sqrt{2}\chi_{\bar{G}F}(\vec{q}\,)&-\chi_{FF}(\vec{q}\,)&-\chi_{\bar{G}G}(\vec{q}\,)\end{bmatrix},$ (36a)
@@ II.2.3 Other thermodynamic potentials
Equations (18)-(21) for the normal state remain valid in the superconducting phase except Eq. (19a) for the internal energy ${\cal E}$, for which the anomalous contribution should be added as
@@ III Numerical Results
Figure 1: Temperature dependences of the Lagrange multiplier $\lambda$ and the first-order anomalous self-energy $\Delta_{1}$ in comparison with that of the mean-field energy gap $\Delta_{\rm MF}$.
We have performed a fully self-consistent numerical calculation of the thermodynamic potentials for the low-density attractive Hubbard model considered previously.Kita24 ; Kita25 Specifically, the band structure is given by $\xi_{\bf k}=k^{2}-\mu$ with $0\leq k\leq 10$ in units of $12m\!=\!k_{\rm F}\!=\!\varepsilon_{\rm F}^{0}=$ 1, where $m$ is the electron mass, $k_{\rm F}$ is the Fermi momentum, and $\varepsilon_{\rm F}^{0}=\mu_{0}(T=0)$ denotes the non-interacting Fermi energy. Temperature dependences of basic parameters were obtained as Fig. 1 at an intermediate coupling of $UN(0)=-0.09$,Kita24 where $\Delta_{\rm MF}$ and $T_{{\rm c}0}\!=\!0.1767$ are the mean-field results on the energy gap and transition temperature, $\lambda$ is the prefactor of the 1PR contribution to the self-energy in Eq. (29) of Ref. Kita25, , and $\Delta_{1}$ is the energy gap evaluated by Eq. (28b) in terms of the fully self-consistent $F$. The parameters $\lambda$ and $\Delta_{1}$ start to develop at $T_{\rm 1PR}=0.100$ and $T_{\rm c}=0.0654$, respectively. Whereas $T_{\rm c}$ signifies the superconducting transition temperature, $T_{\rm 1PR}$ is not directly connected with any broken symmetry but only characterized by the emergence of the 1PR structure in the self-energy.Kita24 As shown in a previous study,Kita25 this $T_{\rm 1PR}$ marks the onset of a pseudogap behavior in the single-particle density of states; it may be denoted more appropriately as $T_{\rm pg}$ based on the observable characteristic. Given below are temperature dependences of the basic thermodynamic potentials corresponding to the phase diagram of Fig. 1.

### v2

@@ I Introduction
Quantum field theory has proved extremely useful for describing properties of many-particle systems, as can be seen by looking into classic textbooks.KB62 ; AGD63 ; FW71 ; NO88 ; Mahan00 However, the main focus there is on Green’s functions that are relevant to one- and two-particle properties rather than thermodynamic potentials. Moreover, ordered phases such as superconductivity are treated almost exclusively within the mean-field framework. Indeed, there seems to have been few detailed studies on how to calculate thermodynamic potentials of normal and ordered phases with correlations microscopically and continuously through phase transitions. One of the exceptional ones was given by Rainer and SereneRS76 ; SR83 on thermodynamics of superfluid 3He, where correlation effects are treated perturbatively as strong-coupling effects on the mean-field results by using the LW functional. Another study to be mentioned was given by Haussmann et alHRCZ07 on thermodynamics of the BCS-BEC crossover of atomic gases based on the attractive Hubbard model, who adopted the De Dominicis-Martin (DDM) formalismDDM64 of expressing $\Omega$ as a functional of Green’s functions and two-particle vertices in addition to the LW framework. However, (i) their mixed use of the DDM and LW formalisms and (ii) the renormalization of the bare interaction vertex obscure the exact procedure to obtain the thermodynamic potentials.
[…]
@@ II.1.1 Grand thermodynamic potential
$\,\times\frac{\delta\Sigma(\vec{k}^{\prime})}{\delta G(\vec{k})}e^{i\varepsilon_{n^{\prime}}0_{+}}-2T\Sigma(\vec{k})+\frac{\delta\Phi}{\delta G(\vec{k})}.$ (11)
@@ II.1.2 $\Phi$ functional
which holds for the second-order contribution to $\Phi$ in the normal state. However, the former expression with $\chi_{GG}(\vec{q})$ is far more suitable numerically because of the inequality $|\chi_{GG}(\vec{q})|\ll|\chi_{G\bar{G}}(\vec{q})|$ that holds for the attractive Hubbard model at low temperatures. Indeed, Eq. (17) indicates that the tiny second-order contribution given in terms of the particle-hole bubble $\chi_{GG}(\vec{q})$ results by the expression in terms of the particle-particle bubble $\chi_{G\bar{G}}(\vec{q})$ through the cancelation of its huge real and imaginary numbers. Hence, we have calculated the normal-state $\Phi$ by Eq. (14) above.
@@ II.1.3 Other thermodynamic potentials
[…]
@@
[…]
@@
[…]
@@ II.2.2 $\Phi$ functional
$\underline{\chi}^{({\rm c})}(\vec{q}\,)\equiv\begin{bmatrix}\vskip 2.84526pt\chi_{-}(\vec{q}\,)&\sqrt{2}\chi_{GF}(\vec{q}\,)&-\sqrt{2}\chi_{\bar{G}F}(\vec{q}\,)\\ \vskip 2.84526pt\sqrt{2}\chi_{GF}(\vec{q}\,)&-\chi_{G\bar{G}}(\vec{q}\,)&-\chi_{FF}(\vec{q}\,)\\ -\sqrt{2}\chi_{\bar{G}F}(\vec{q}\,)&-\chi_{FF}(\vec{q}\,)&-\chi_{\bar{G}G}(\vec{q}\,)\end{bmatrix}$, (36a)
@@ II.2.3 Other thermodynamic potentials
[…]
@@ III Numerical Results
Figure 1: Temperature dependences of the Lagrange multiplier $\lambda$, the first-order anomalous self-energy $\Delta_{1}$ in comparison with the mean-field energy gap $\Delta_{\rm MF}$, and the single-particle density of states at zero energy $D(0)$.
We have performed a fully self-consistent numerical calculation of the thermodynamic potentials for the low-density attractive Hubbard model considered previously.Kita24 ; Kita25 Specifically, the band structure is given by $\xi_{\bm{k}}=k^{2}-\mu$ with $0\leq k\leq 10$ in units of $2m\!=\!k_{\rm F}\!=\!\varepsilon_{\rm F}^{0}=1$, where $m$ is the electron mass, $k_{\rm F}$ is the Fermi momentum, and $\varepsilon_{\rm F}^{0}=\mu_{0}(T=0)$ denotes the non-interacting Fermi energy. Temperature dependences of basic parameters were obtained as Fig. 1 at an intermediate coupling of UN(0)=0.09UN(0)=-0.09,Kita24 where $\Delta_{\rm MF}$ and $T_{{\rm c}0}\!=\!0.1767$ are the mean-field results on the energy gap and transition temperature, $\lambda$ is the prefactor of the 1PR contribution to the self-energy in Eq. (29) of Ref. Kita25, , $\Delta_{1}$ is the energy gap evaluated by Eq. (28b) in terms of the fully self-consistent $F$, and $D(0)$ denotes the single-particle density of states at zero energy defined by Eq. (40) of Ref. Kita25, . The parameters $\lambda$ and $\Delta_{1}$ start to develop at $T_{\rm 1PR}=0.100$ and $T_{\rm c}=0.0654$, respectively. Whereas $T_{\rm c}$ signifies the superconducting transition temperature, $T_{\rm 1PR}$ is not directly connected with any broken symmetry but only characterized by the emergence of the 1PR structure in the self-energy.Kita24 As shown in a previous study,Kita25 this $T_{\rm 1PR}$ marks the onset of a pseudogap behavior in the single-particle density of states, as seen in a sharp drop of $D(0)$ at $T_{\rm 1PR}$ in Fig. 1; it may be denoted more appropriately as $T_{\rm pg}$ based on the observable characteristic. Given below are temperature dependences of the basic thermodynamic potentials corresponding to the phase diagram of Fig. 1.